# Are We Truly Forgetting? A Critical Re-examination of Machine Unlearning Evaluation Protocols

### v2

@@ Abstract
Machine unlearning is a process to remove specific data points from a trained model while maintaining the performance on retain data, addressing privacy or legal requirements. Despite its importance, existing unlearning evaluations tend to focus on logit-based metrics (i.e., accuracy) under small-scale scenarios. We observe that this could lead to a false sense of security in unlearning approaches under real-world scenarios. In this paper, we conduct a new comprehensive evaluation that employs representation-based evaluations of the unlearned model under large-scale scenarios to verify whether the unlearning approaches genuinely eliminate the targeted forget data from the model’s representation perspective. Our analysis reveals that current state-of-the-art unlearning approaches either completely degrade the representational quality of the unlearned model or merely modify the classifier (i.e., the last layer), thereby achieving superior logit-based evaluation metrics while maintaining significant representational similarity to the original model. Furthermore, we introduce a rigorous unlearning evaluation setup, in which the forgetting classes exhibit semantic similarity to downstream task classes, necessitating that feature representations diverge significantly from those of the original model, thus enabling a more rigorous evaluation from a representation perspective. We hope our benchmark serves as a standardized protocol for evaluating unlearning algorithms under realistic conditions.
[…]
@@ 1 Introduction
Neural networks have driven transformative advancements across a multitude of domains, such as image classification [10, 18, 25]. A key contributor to this progress has been the substantial enhancement of the model capability, facilitated by the training of sophisticated models on vast datasets [1]. However, these successes have also raised significant privacy concerns about these datasets, particularly related to the unchecked use of sensitive data [4]. As privacy rights, including the ‘right to be forgotten’ (e.g., the GDPR legislation [3, 27]), have gained increasing recognition, the need for effective solutions has become urgent. As a result, this has led to the emergence of machine unlearning as a critical area of research [28].
The primary objective of machine unlearning is to effectively eliminate the influence of specific data from a pre-trained model while preserving the knowledge derived from other data [28]. The most straightforward approach, known as exact unlearning, involves retraining the model from scratch using a refined dataset [26] (hereafter referred to as the “retrained model”). While this approach is regarded as the gold standard because it guarantees complete elimination of any influence from data designated for removal, it incurs significant computational expenses due to the necessity of retraining models on refined datasets whenever unlearning is required. To overcome these limitations, several approximate unlearning algorithms have been proposed, offering more efficient alternatives while striving to maintain the goal of machine unlearning [16]. In image classification tasks, approximate unlearning algorithms are typically evaluated based on their ability to produce models that achieve classification accuracy similar to that of the exact unlearning on the given dataset (e.g., retain and forget sets) [28], as illustrated on the left side of Fig. 1. Following this evaluation method, several methods have induced unlearning through techniques such as using Random Labeling [13], Gradient Ascent, [32] and metric learning [2], while mitigating the over-forgetting by incorporating additional regularization terms [23, 12, 5, 2].
@@
Despite recent advancements, current evaluations of approximate unlearning research remain confined to logit-based evaluation (e.g., classification accuracy) on small-scale unlearning tasks and datasets, such as unlearning a limited number of classes for models trained on CIFAR-10 and CIFAR-100 [22], as shown in Fig. 1(a). Such narrow evaluation scopes raise significant concerns about the scalability of these methods to larger datasets and real-world unlearning scenarios. Privacy considerations assume heightened significance in larger model architectures (e.g., foundation models), as these systems are trained on extensive datasets frequently sourced from web-crawled images, subsequently employed across diverse downstream applications. This limitation overlooks the broader implications of unlearning algorithms on the model’s foundational component (i.e., encoder), potentially leading to incomplete or misleading assessments of its effectiveness.
@@
In this paper, we propose a comprehensive new evaluation framework that addresses the limitations of existing small-scale, logit-based evaluations by incorporating assessments of unlearned models ($\theta_{u}$) from the perspective of feature representations in large-scale settings. As illustrated in Fig. 1(b), our framework evaluates the representations of unlearned models using Centered Kernel Alignment (CKA) [20] to compare the feature representations of the unlearned models with those of a model trained only on the retain set ($\theta_{r}$). Additionally, we apply $k$-Nearest Neighbors ($k$-NN) analysis on the feature space across various downstream datasets to assess the quality of the unlearned model’s feature representations and compare the results with those of $\theta_{r}$. Through these evaluation methods, we re-evaluate multiple state-of-the-art unlearning algorithms in the large-scale unlearning scenario. Fig. 2 illustrates key findings from experiments on unlearning 100 random classes in ImageNet-1K using ResNet-50: while logit-based evaluation (e.g., classification accuracy) demonstrates successful unlearning by showing a decreasing gap between the unlearned models ($\theta_{u}$) and the retrained model ($\theta_{r}$) (Fig. 2, left), a representation-based evaluation using CKA reveals a critical limitation (Fig. 2, right). In an ideal scenario, where unlearning is truly effective, the unlearned model should be indistinguishable from the retrained model, with the ideal value being 10. However, as shown in Fig. 2 (right), all unlearned models exhibit greater similarity to the original model than to the retrained model. This highlights a fundamental issue: existing unlearning methods fail to sufficiently erase learned information at the representation level. This reinforces the need for representation-based evaluation to rigorously assess whether unlearning truly eliminates the influence of the targeted data.
Furthermore, random class forgetting has been widely adopted as a standard evaluation scenario in unlearning studies [5, 12, 2, 23]. However, we highlight a critical limitation of this approach: the forgetting process may be inadequately assessed, as similar representational knowledge can persist across multiple alternative classes, making it difficult to verify whether the targeted data has been truly unlearned. To address this issue, we propose a novel unlearning evaluation paradigm, Top Class-wise Forgetting, in which the classes selected for unlearning are semantically similar to those in downstream tasks. This setup requires a substantial divergence of feature representations from the original model, offering a more rigorous assessment of the unlearning algorithm’s effectiveness. We then evaluate existing unlearning methods using both logit-based and representation-based metrics across diverse downstream datasets. Our goal is to establish a reusable and scalable benchmark for the unlearning community, enabling standardized and holistic evaluation of future methods. In summary, the core contributions are as follows:
  • We identify fundamental limitations in current unlearning evaluations, which overly rely on logit-based metrics under small-scale settings.
  • To address this, we propose a unified benchmark framework that incorporates representation-based metrics and reveals that many unlearning methods fail to alter internal representations.
@@ 2 Related Works
Machine unlearning in image classification. Early research on machine unlearning in neural networks has primarily focused on image classification, leading to the development of various unlearning algorithms with distinct strategies [28]. A widely used approach, Gradient Ascent (GA), maximizes the loss on the forget set to erase its influence, making it one of the most fundamental yet widely adopted techniques in recent studies [14, 31, 15, 37, 5]. Similarly, Fine-tuning achieves unlearning by inducing catastrophic forgetting on forget set, updating the model solely on retain set [14, 36]. More sophisticated unlearning techniques have been developed: SCRUB [23] employs a knowledge distillation framework to selectively forget knowledge from the forget set while retaining other knowledge. L2UL [5] enhances unlearning by incorporating two regularization terms, which mitigate forgetting at both the feature and weight levels when combined with GA. SalUn [12] modifies the most influential parameters through weight saliency, enabling selective unlearning. SCAR [2] leverages metric learning alongside a distillation-based strategy, aligning forget set representations to the nearest incorrect class using the Mahalanobis distance.
@@
Evaluation and analysis of representations. Transfer learning has become a foundational technique in machine learning, facilitating the adaptation of pre-trained models to diverse downstream tasks [29]. Beyond its practical applications, it is widely used to assess the quality of the learned representations. For instance, it has been employed to compare the learned representations of different model architectures [21] and algorithms across various domains, such as self-supervised learning [8, 17] and continual learning [6]. This evaluation is typically performed using metrics such as $k$-Nearest Neighbors ($k$-NN), which measure how well the learned representations generalize in various downstream tasks. Centered Kernel Alignment (CKA) [20] has been widely used to compare feature representations across models, providing insights into their structural similarities and differences. Furthermore, t-SNE [33] has been extensively utilized to visualize feature representations, offering an intuitive means to compare and analyze how different models organize learned representations.
@@
In contrast to prior work, we revisit the foundations of unlearning evaluation by exposing key limitations in conventional protocols and propose a unified, representation-aware benchmark that enables more realistic and rigorous assessment.
@@ 3 Why We Need Better Benchmarks: An Empirical Study
In this section, we present a comprehensive analysis of existing machine unlearning methods, focusing on both logit-based and representation-based evaluations in the large-scale setup. We review the classical unlearning evaluation protocols and introduce a new evaluation metric and setup in Sec. 3.1. Then, we introduce the motivation of our approach by highlighting limitations in current evaluation frameworks in Sec. 3.2, introduce our proposed evaluation metric in Sec. 4, present experimental findings under large-scale settings from a representation perspective in Sec. 4.1, and show results under top class-wise forgetting scenarios from a transfer learning perspective in Sec. 4.2.
3.1 Problem Setup and Evaluation Metrics for Unlearning
@@
Problem setup and objective of machine unlearning in the classification task.   We define the entire dataset as $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}$ represents the input image, $y_{i}$ denotes its corresponding label, and $N$ is the total number of samples in the dataset. The subset of data targeted for unlearning, referred to as the forget set, is denoted by $\mathcal{D}_{f}\subset\mathcal{D}$, and consists of samples from specific classes that the model is intended to forget. Retain set, denoted by $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$, includes all samples in $\mathcal{D}$ that are not part of the forget set. Additionally, we introduce the test forget set, denoted as $\mathcal{D}^{\text{te}}_{f}$, which consists of data points from the same classes as $\mathcal{D}_{f}$ but not included in the train set. Similarly, the test retain set, denoted as $\mathcal{D}^{\text{te}}_{r}$, contains data points from the same classes as $\mathcal{D}_{r}$. We assume that the original model $\theta$ is trained on the entire dataset $\mathcal{D}$ and subsequently unlearns information related to the forget set $\mathcal{D}_{f}$. Note that the goal of machine unlearning is to produce an unlearned model $\theta_{u}$ that retains the knowledge of $\mathcal{D}_{r}$ while achieving the unlearning of $\mathcal{D}_{f}$. Additionally, we consider a retrained model $\theta_{r}$, which is trained on $\mathcal{D}_{r}$ from scratch, as an oracle for comparison in the unlearning process.

### v3

@@ Abstract
Machine unlearning is a process to remove specific data points from a trained model while maintaining the performance on the retain data, addressing privacy or legal requirements. Despite its importance, existing unlearning evaluations tend to focus on logit-based metrics under small-scale scenarios. We observe that this could lead to a false sense of security in unlearning approaches under real-world scenarios. In this paper, we conduct a comprehensive evaluation that employs representation-based evaluations of the unlearned model under large-scale scenarios to verify whether the unlearning approaches truly eliminate the targeted data from the model’s representation perspective. Our analysis reveals that current state-of-the-art unlearning approaches either completely degrade the representational quality of the unlearned model or merely modify the classifier, thereby achieving superior logit-based performance while maintaining representational similarity to the original model. Furthermore, we introduce a novel unlearning evaluation scenario in which the forgetting classes exhibit semantic similarity to downstream task classes, necessitating that feature representations diverge significantly from those of the original model, thus enabling a more thorough evaluation from a representation perspective. We hope our benchmark will serve as a standardized protocol for evaluating unlearning algorithms under realistic conditions.
keywords:
Machine Unlearning, Unlearning Evaluation Benchmark, Representation Learning, Data Privacy, Transfer Learning, Machine Learning
journal: Engineering Applications of Artificial Intelligence
@@ 1 Introduction
Neural networks have driven transformative advancements across a multitude of domains, such as image classification [14, 21, 32]. A key contributor to this progress has been the substantial enhancement of the model’s capabilities, facilitated by the training of sophisticated models on vast datasets [2]. However, these successes have also raised significant privacy concerns about these datasets, particularly related to the unchecked use of sensitive data [5]. As privacy rights, including the ‘right to be forgotten’ (e.g., the GDPR legislation [4, 35]), have gained increasing recognition, the need for effective solutions has become urgent. As a result, this has led to the emergence of machine unlearning as a critical research area [36].
The primary objective of machine unlearning is to effectively eliminate the influence of specific data from a pre-trained model while preserving the knowledge derived from other data [36]. The most straightforward approach, known as exact unlearning, involves retraining the model from scratch using a refined dataset [34] (hereafter referred to as the “retrained model”). This approach is regarded as the gold standard as it guarantees the complete elimination of the targeted data. However, it is computationally expensive, requiring retraining for each unlearning request. To overcome these limitations, several approximate unlearning algorithms have been proposed. They offer more efficient alternatives while maintaining the goal of machine unlearning. In image classification tasks, approximate unlearning algorithms are typically evaluated based on their ability to produce models that achieve classification accuracy similar to that of the retrained model on the given dataset (e.g., retain and forget sets) [36], as illustrated on the left side of Fig. 1. Following this evaluation method, several techniques have induced unlearning through methods such as Gradient Ascent [44], Random Labeling [18], and metric learning [3], while mitigating over-forgetting by incorporating additional regularization terms [28, 15, 6, 3].
@@
Despite recent advancements, current unlearning evaluations remain narrow. They primarily focus on logit-based metrics (e.g., accuracy, Membership Inference Attack) and small-scale tasks, such as unlearning one class or a few classes in CIFAR-10, as shown in Fig. 1 and Table 1. Such narrow evaluation scopes raise significant concerns about the scalability of these methods to larger datasets and real-world scenarios. Privacy considerations assume heightened significance in larger model architectures (e.g., foundation models), as these systems are trained on extensive datasets, frequently sourced from web-crawled images, and are subsequently employed across diverse downstream applications. This limitation overlooks the broader implications of unlearning algorithms on the model’s foundational component (i.e., encoder), potentially leading to incomplete assessments of its effectiveness.
Table 1: Comparison of evaluation protocols in recent machine unlearning literature. While current unlearning methods are evaluated on small-scale datasets for unlearning, our work utilizes ImageNet-1K for unlearning, which consists of about 1.28 million images.
Work Backbone Model Unlearning Dataset Unlearning Classes Downstream Dataset
[44] ResNet-18 CIFAR-10/100 (#60K) -
[18] ResNet-18 MNIST (#70K), etc. 1
[8] ResNet-18, VGG-11, AllCNN SVHN (#99K), etc. 1-10
[15] ResNet-18 CIFAR-10 (#60K) 1
[12] ResNet-18, VGG Tiny ImageNet (#110K), etc. 1
[52] ResNet-18/34/50/101 SVHN (#99K), etc. 1
[28] ResNet-18, AllCNN CIFAR-10 (#60K), etc. 1
[3] ResNet-18/34/50, AllCNN, ViT-B16 Tiny ImageNet (#110K), etc. 1
Ours ResNet-50, Swin-T, ConvNeXt ImageNet-1K (#1.28M) 100-300
In this paper, we propose a holistic unlearning evaluation framework that addresses the limitations of small-scale, logit-based evaluations. We incorporate assessments of unlearned models from the perspective of feature representations in large-scale settings. As illustrated in Fig. 1, our framework evaluates the representations of the unlearned models using Centered Kernel Alignment (CKA) [26] to compare the feature representations of the unlearned models with those of the retrained model ($\theta_{r}$). Additionally, we apply $k$-Nearest Neighbors ($k$-NN) analysis on the feature space across various downstream datasets to assess the quality of the unlearned model’s feature representations and compare the results with those of the retrained model. Through these evaluation methods, we re-evaluate multiple state-of-the-art unlearning algorithms in large-scale scenarios. Fig. 2 illustrates key findings from experiments on unlearning 100 random classes in ImageNet-1K using ResNet-50. While logit-based evaluation (e.g., classification accuracy) demonstrates successful unlearning by showing a decreasing gap between the unlearned models ($\theta_{u}$) and the retrained models ($\theta_{r}$) (Fig. 2, left), a representation-based evaluation using CKA reveals a critical limitation. If unlearning is truly effective, the unlearned model should be indistinguishable from the retrained model, with the ideal value being 10. However, as shown in Fig. 2 (right), all unlearned models exhibit greater similarity to the original model than to the retrained model. Furthermore, the order of effectiveness of these algorithms varies depending on whether a logit-based or representation-based evaluation is used. These findings highlight a fundamental issue: existing unlearning methods fail to sufficiently erase learned information at the representation level, and they were not evaluated from multifaceted perspectives. These findings reinforce the need for representation-based evaluation to rigorously assess whether unlearning truly eliminates the influence of the targeted data.
@@
Furthermore, random class forgetting has been widely adopted as a standard evaluation scenario in unlearning studies [6, 15, 3, 28]. However, we highlight a critical limitation of this approach: the forgetting process may be inadequately assessed, as similar representational knowledge can persist across multiple alternative classes, making it difficult to verify whether the targeted data has truly been unlearned. To address this issue, we propose a novel unlearning evaluation paradigm, Top Class-wise Forgetting, in which the classes selected for unlearning are semantically similar to those in downstream tasks. This setup requires a substantial divergence of feature representations from the original model, offering a more rigorous assessment of the unlearning algorithm’s effectiveness. We then evaluate existing unlearning methods using both logit-based and representation-based metrics across diverse downstream datasets. Our goal is to establish a rigorous benchmark for the community, enabling a holistic evaluation of future methods. In summary, our contributions are as follows:
  • 1.
    We identify fundamental limitations in current evaluations, which rely on logit-based metrics under small-scale scenarios.
  • 2.
    To address this, we propose a unified benchmark that incorporates representation-based metrics and reveals that existing methods fail to alter internal representations. Furthermore, we conduct various experiments in large-scale settings (i.e., evaluations on ImageNet-1K using ResNet-50, Swin-T, and ConvNeXt) to determine whether current baselines can truly unlearn the model’s knowledge.
  • 3.
@@ 2 Related Works
Machine unlearning in image classification. Early research on machine unlearning in neural networks has primarily focused on image classification, leading to the development of various unlearning algorithms with distinct strategies [36]. A widely used approach, Gradient Ascent (GA), maximizes the loss on the forget set to erase its influence, making it one of the most fundamental yet widely adopted techniques in recent studies [18, 43, 19, 50, 6]. Similarly, Fine-tuning achieves unlearning by inducing catastrophic forgetting on the forget set, updating the model solely on the retain set [18, 48]. More sophisticated unlearning techniques have been developed: SCRUB [28] employs a knowledge distillation framework to selectively forget knowledge from the forget set while retaining other knowledge. L2UL [6] enhances unlearning by incorporating two regularization terms that mitigate forgetting at both the feature and weight levels when combined with GA. SalUn [15] modifies the most influential parameters through weight saliency, enabling selective unlearning. SCAR [3] leverages metric learning alongside a distillation-based strategy, aligning the forget set representations to the nearest incorrect class using the Mahalanobis distance. Beyond image classification, recent works have begun to address the complexities of unlearning in generative models, specifically those pre-trained on large-scale datasets such as ImageNet-1K [30, 16].
@@
Evaluation and analysis of representations. Transfer learning has become a foundational technique in machine learning, facilitating the adaptation of pre-trained models to diverse downstream tasks [37]. Beyond its practical applications, it is widely used to assess the quality of the learned representations. For instance, it has been employed to compare the learned representations of different model architectures [27] and algorithms across various domains, such as self-supervised learning [10, 20] and continual learning [7]. This evaluation is typically performed using metrics such as $k$-Nearest Neighbors ($k$-NN), which measures how well the learned representations generalize in various downstream tasks. Centered Kernel Alignment (CKA) [26] has been widely used to compare feature representations across models, providing insights into their structural similarities and differences. Furthermore, t-SNE [45] has been extensively utilized to visualize feature representations, offering an intuitive means to compare and analyze how different models organize learned representations.
@@
In contrast to prior work, we revisit the foundations of unlearning evaluation by exposing key limitations in conventional protocols and proposing a unified, representation-aware benchmark that enables a more realistic and rigorous assessment.
Traditional evaluation framework in machine unlearning. In the literature on machine unlearning, particularly for image classification tasks, logit-based evaluation metrics have primarily been used to quantify the effectiveness of unlearning algorithms on small-scale scenarios [18, 44, 12, 28, 9, 23, 15, 6, 3, 8, 52, 53]. The most common metrics are Forget Accuracy and Retain Accuracy. Forget Accuracy measures how effectively the model removes knowledge related to the data intended to be forgotten, typically by evaluating the model’s performance on the forget set. In contrast, Retain Accuracy assesses how well the model preserves its performance on the retain set, thus ensuring that unlearning does not degrade the model’s utility on the retain set. Additionally, Membership Inference Attacks (MIA) have become a standard evaluation metric in machine unlearning to quantify the privacy risks associated with residual information about the forget data in models [5, 42, 17, 11]. Introduced in early privacy research [40], MIA simulates an adversary’s attempt to determine whether a given data sample was included in the training set by analyzing the model’s outputs. In the context of machine unlearning, MIA is used to assess the effectiveness of unlearning algorithms by measuring whether an attacker can still infer membership in the forget set after the unlearning process. Recent unlearning work [23] employs MIA to rigorously evaluate their unlearning method, demonstrating that their method significantly reduces the success rate of MIA on the forget set compared to prior methods. While these metrics have served as fundamental tools for evaluation, they predominantly focus on output-level behaviors and may not fully capture changes in the internal representation or feature space of the model. Consequently, recent works have begun to emphasize the need for representation-based evaluation frameworks [22, 29, 41, 39].
Representation-based evaluation in machine unlearning. Recent research in machine unlearning has begun to acknowledge the limitations of conventional logit-based evaluation, shifting towards methods that analyze the model’s internal representations [22, 29, 41, 39]. [22] proposed an approach based on information theory. They introduced the Information Difference Index (IDI), a white-box metric that quantifies residual information by measuring the mutual information between intermediate features and the forgetting labels. IDI is a powerful tool for detecting subtle information leakage that black-box metrics might miss by tracking the information within the model. [29] introduced the concept of Knowledge Deletion (KD), pointing out that existing unlearning methods often modify only the classification head while leaving the feature extractor largely unchanged. To cope with this, they proposed the Knowledge Retention (KR) score, a metric that measures the accuracy of a linear classifier retrained on the frozen feature extractor of an unlearned model.
In contrast to these prior works, our evaluation framework offers two key advantages. First, it does not require any additional training or label supervision, enabling a training-free and computationally efficient evaluation that can be applied across diverse downstream datasets (e.g., Office-Home). Second, unlike IDI or KR, which still rely on accuracy (e.g., linear head retraining with binary labels in IDI or classification in KR), our approach directly measures the representational distance between the unlearned model, the original model, and the gold standard, allowing us to quantitatively assess how unlearning impacts the final model. Furthermore, our framework goes beyond proposing a representation-based metric. It also introduces large-scale evaluation settings, practical transferability assessment across diverse downstream datasets, and the Top Class-wise Forgetting scenario. While these prior works evaluate small-scale scenarios such as CIFAR-10, Tiny ImageNet, or ImageNet-1K with only 5-class forgetting, our framework targets large-scale unlearning involving 100–300 classes and proposes the Top Class-wise Forgetting setup to address the challenges of evaluating unlearning performance under random class forgetting. In summary, our evaluation framework advances beyond previous evaluation metrics by incorporating the training-free, computationally efficient evaluation metric from a transfer learning perspective and a novel, large-scale scenario.
@@ 3 Why We Need Better Benchmarks: An Empirical Study
In this section, we formulate the problem and present a comprehensive analysis of existing machine unlearning methods, focusing on both logit-based and representation-based evaluations in large-scale settings. We review classical unlearning evaluation protocols and introduce novel evaluation metrics and scenarios. Next, we motivate our approach by highlighting the limitations of current evaluation frameworks. Then, in Sec. 4, we introduce our proposed evaluation metric, present experimental results under large-scale settings from a representation perspective, and report findings from Top Class-wise Forgetting scenarios from a transfer learning perspective.
3.1 Problem Setup
@@
We define the entire dataset as $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, where $x_{i}$ represents the input image, $y_{i}$ denotes its corresponding label, and $N$ is the total number of samples in the dataset. The subset of data targeted for unlearning, referred to as the forget set, is denoted by $\mathcal{D}_{f}\subset\mathcal{D}$. The retain set, denoted by $\mathcal{D}_{r}=\mathcal{D}\setminus\mathcal{D}_{f}$, includes all samples in $\mathcal{D}$ that are not part of the forget set. Additionally, we define the test forget set, denoted as $\mathcal{D}^{\text{te}}_{f}$, which consists of data points from the same classes as $\mathcal{D}_{f}$ but are not included in the train set. Similarly, the test retain set, denoted as $\mathcal{D}^{\text{te}}_{r}$, contains data points from the same classes as $\mathcal{D}_{r}$. We assume that the original model $\theta_{o}$ is trained on the entire dataset $\mathcal{D}$ and subsequently unlearns the knowledge related to the forget set $\mathcal{D}_{f}$. Note that the goal of machine unlearning is to produce an unlearned model $\theta_{u}$ that retains the knowledge of $\mathcal{D}_{r}$ while achieving the unlearning of $\mathcal{D}_{f}$. Additionally, we consider a retrained model $\theta_{r}$, which is trained on $\mathcal{D}_{r}$ from scratch, as an oracle for comparison in the unlearning process.
3.2 Logit-based Evaluation Metrics